### v1

# Millinovae: a new class of transient supersoft X-ray sources without a nova eruption

Abstract
Some accreting binary systems containing a white dwarf (such as classical novae or persistent supersoft sources) are seen to emit low energy X-rays with temperatures of 106 K and luminosities exceeding erg s-1. These X-rays are thought to originate from nuclear burning on the white dwarf surface, either caused by a thermonuclear runaway (classical novae) or a high mass accretion rate that sustains steady nuclear burning (persistent sources). The discovery of transient supersoft X-rays from ASASSN-16oh challenged these ideas, as no signatures of nuclear fusion were detected, and the origin of these X-rays remains controversial. It was unclear whether this star was one of a kind or representative of a larger, as yet undiscovered, group. Here we present the discovery of 29 stars located in the direction of the Magellanic Clouds exhibiting long-duration, symmetrical optical outbursts similar to that seen in ASASSN-16oh. We observed one of these objects during an optical outburst and found it to be emitting transient supersoft X-rays, while no signatures of mass ejection (indicative of a classical nova eruption) were detected. We therefore propose that these objects form a homogeneous group of transient supersoft X-ray sources, which we dub ‘millinovae’ because their optical luminosities are approximately a thousand times fainter than those of ordinary classical novae.
1 Introduction
A classical nova eruption occurs on the surface of a white dwarf that is accreting matter from its non-degenerate companion via an accretion disk. A layer of fresh, usually hydrogen-rich gas accumulates on the white dwarf surface and becomes degenerate. The gas is heated, and when hydrogen ultimately ignites, a thermonuclear runaway ejects about of matter from the system with velocities from hundreds to thousands of km s-1(Bode & Evans, 2008). As the ejecta expands and becomes optically thin, the exposed white dwarf surface is observed to emit supersoft X-rays due to stable hydrogen burning of the leftover gas (Osborne, 2015; Page & Shaw, 2022). The maximum X-ray luminosities of classical novae are typically erg s-1 (e.g., Henze et al., 2014). In the optical passbands, they normally reach absolute magnitudes from to (Bode & Evans, 2008).
The accretion disk surrounding the white dwarf is subject to a thermal instability that can lead to dwarf nova outbursts. They occur when gas in the disk reaches a critical temperature, and its viscosity abruptly changes, increasing the mass accretion rate onto the white dwarf (Lasota, 2001). If the instability develops in the outer regions of the accretion disk (outside-in outbursts), the rise to maximum brightness is rapid. However, the light curves of inside-out outbursts (which start in the inner disk regions) are more symmetric, with a relatively slow rise to the peak (Lasota, 2001). Dwarf novae in outburst reach optical absolute magnitudes from to , depending on their orbital period (longer orbital periods mean larger accretion disks, and hence become brighter; Patterson 2011). The X-ray luminosities of dwarf novae in outburst are typically erg s-1, about six orders of magnitude lower than those of classical novae (Schwope et al., 2024; Rodriguez et al., 2024).
Therefore, the discovery of supersoft X-ray emission from the optical transient ASASSN-16oh (Maccarone et al., 2019) was a surprise. The X-ray emission was consistent with that of a 900,000 K blackbody and a luminosity of about erg s-1, similar to that seen from persistent supersoft X-ray sources and classical novae. However, the optical properties of the transient (light curve shape, its amplitude, peak absolute magnitude, optical spectra) did not match those of classical novae at all.
ASASSN-16oh was first detected in December 2016 by the All-Sky Automated Survey for Supernovae (ASAS-SN; Shappee et al. 2014) as a transient in the Small Magellanic Cloud (SMC). Its archival light curve (upper panel of Figure 1) collected by the Optical Gravitational Lensing Experiment (OGLE; Udalski et al. 2015) revealed that the optical outburst commenced at least four months before the ASAS-SN discovery. The object exhibited irregular variability before (and after) the outburst, with a mean brightness of about and . It peaked at , which corresponds to an absolute magnitude of , assuming the SMC distance modulus of (Graczyk et al., 2020) and reddening of (Skowron et al., 2021). The full optical light curve does not match those of classical novae, which normally rise to maximum within hours to a few days (not months). Additionally, the optical spectra revealed narrow ( km s-1) Balmer and He ii emission lines (Maccarone et al., 2019), unlike the broad features (several hundred to a few thousand km s-1) usually observed in classical novae.
The unusual outburst light curve, narrow optical emission lines, and X-ray spectrum led Maccarone et al. (2019) to propose that the X-rays come from a spreading layer—a belt around the white dwarf’s equator near the inner disk edge in which a large fraction of the total accretion energy is emitted. However, the nature of ASASSN-16oh remains controversial, with alternative models being based on a non-ejecting nova outburst (Hillman et al., 2019; Kato et al., 2020).
Given the unusual and unexplained optical and X-ray properties of ASASSN-16oh, we decided to investigate whether this star is one of a kind or representative of a larger, as yet undiscovered, group. We searched for objects with outburst properties similar to those seen in ASASSN-16oh among the 20-year-long light curves of 76 million stars observed toward the Magellanic Clouds by the OGLE survey (Udalski et al., 2015).
2 OGLE Photometric Data
The OGLE photometric data have been collected using the 1.3-m Warsaw Telescope located at Las Campanas Observatory, Chile, during the OGLE-III (2001–2009; Udalski 2003) and OGLE-IV (2010–2024; Udalski et al. 2015) phases of the project. Different instruments were used in different phases of the survey. The OGLE-III camera consisted of eight CCD detectors, with a plate scale of arcsec pixel-1. The OGLE-IV camera had the same pixel scale but comprised thirty-two detectors, providing a field of view of 1.4 deg2. The typical exposure time was 150 s. The photometric data were extracted using a custom version of the Difference Image Analysis algorithm (Tomaney & Crotts, 1996; Alard & Lupton, 1998) implemented by Woźniak (2000). The OGLE-III and OGLE-IV data have been reduced using the same reference image (Mróz et al., 2024), enabling us to obtain homogeneous, long-term light curves of all analyzed objects. Photometry was calibrated to the standard - and -band systems. See Udalski (2003) and Udalski et al. (2015) for a detailed description of the observing setup and data calibration.
3 Results
We started with a list of 72,303 objects (54,996 in the Large Magellanic Cloud (LMC) and 17,307 in the SMC) whose light curves contain at least five consecutive data points magnified with respect to the remaining light curve (see Mróz et al. (2024) for more details about the selection of outbursting events). We required that the light curve contains at least one outburst with an amplitude larger than 1 mag and duration between 10 and 600 days. Subsequently, we cross-matched selected objects with the Gaia Data Release 3 archive (Gaia Collaboration et al., 2023) and removed artifacts due to high-proper-motion stars (those with proper motions 10 mas yr-1), which were the largest source of contamination. The light curves of the remaining objects (7412 in the LMC and 2689 in the SMC) were visually inspected, and we selected the initial sample of outbursts similar to those of ASASSN-16oh.
We vetted all selected objects using additional data, including multi-color photometry, Gaia parallaxes and proper motions, and sky images. We removed objects that could be classified as classical novae, supernovae, and active galactic nuclei, leaving us with a list of 29 objects (22 in the LMC and seven in the SMC). Example light curves are presented in Figure 1 (see also Figures 7 and 8 in Appendix A). All objects show long-duration (weeks to months), symmetrical (triangle-shaped) outbursts with amplitudes in ranging from 1.0 to 3.7 mag. Some outbursts are deceivingly symmetric, which led to them being classified as gravitational microlensing events (OGLE-mNOVA-12 = MACHO-LMC-7 and OGLE-mNOVA-13 = MACHO-LMC-23; Alcock et al. 2000), although such events should be achromatic and nonrepeating if they really are due to microlensing. The list of all our selected objects is presented in Table 1.
Given the unique spectral and X-ray characteristics of ASASSN-16oh, we also decided to start the near-real-time monitoring of these selected objects in September 2023, with the goal of triggering follow-up observations should any enter an outburst state.
4 Outburst of OGLE-mNOVA-11
The outburst of OGLE-mNOVA-11 started shortly afterwards, between 2023 October 15.3 and 26.3 and reached a maximum of and on 2023 December 6.1 (Figure 2). This corresponds to and , assuming the LMC distance modulus of (Pietrzyński et al., 2019) and reddening (Skowron et al., 2021). The star returned to quiescence (in ) in late February 2024, about 120 days after the outburst onset. The outburst duration was shorter in .
We obtained a set of low-resolution spectra of OGLE-mNOVA-11 from 2023 November 22 to December 10 (Figure 3) with the Robert Stobie Spectrograph (Kobulnicky et al. 2003; Burgh et al. 2003) mounted on the Southern African Large Telescope (SALT), as part of the SALT Large Science Program on Transients (Appendix B). The spectra cover the range 3500–6700 Å and reveal narrow ( km s-1) emission lines (Balmer lines, He ii 4686 Å). The Bowen blend (a complex of C iii and N iii lines in the range Å) is also clearly detected from 2023 December 2 to 10. The emission lines are redshifted with a mean radial velocity of km s-1, close to that of the LMC systemic value ( km s-1; van der Marel et al. 2002). Both the light curve shape and optical spectra are very different from those of classical novae.
| Name | RA | Decl. | Comments | |||||
| OGLE-mNOVA-01 | 0.059 | ASASSN-16oh | ||||||
| OGLE-mNOVA-02 | 0.042 | |||||||
| OGLE-mNOVA-03 | 0.028 | |||||||
| OGLE-mNOVA-04 | 0.046 | |||||||
| OGLE-mNOVA-05 | 0.016 | candidate | ||||||
| OGLE-mNOVA-06 | 0.011 | |||||||
| OGLE-mNOVA-07 | 0.068 | |||||||
| OGLE-mNOVA-08 | 0.128 | |||||||
| OGLE-mNOVA-09 | 0.122 | candidate | ||||||
| OGLE-mNOVA-10 | 0.119 | |||||||
| OGLE-mNOVA-11 | 0.100 | |||||||
| OGLE-mNOVA-12 | 0.083 | MACHO-LMC-7 | ||||||
| OGLE-mNOVA-13 | 0.124 | MACHO-LMC-23 | ||||||
| OGLE-mNOVA-14 | 0.092 | |||||||
| OGLE-mNOVA-15 | 0.183 | candidate | ||||||
| OGLE-mNOVA-16 | 0.072 | |||||||
| OGLE-mNOVA-17 | 0.105 | |||||||
| OGLE-mNOVA-18 | 0.094 | |||||||
| OGLE-mNOVA-19 | 0.102 | |||||||
| OGLE-mNOVA-20 | 0.078 | |||||||
| OGLE-mNOVA-21 | 0.078 | |||||||
| OGLE-mNOVA-22 | 0.139 | |||||||
| OGLE-mNOVA-23 | 0.099 | |||||||
| OGLE-mNOVA-24 | 0.082 | |||||||
| OGLE-mNOVA-25 | 0.060 | |||||||
| OGLE-mNOVA-26 | 0.124 | |||||||
| OGLE-mNOVA-27 | 0.263 | |||||||
| OGLE-mNOVA-28 | 0.141 | |||||||
| OGLE-mNOVA-29 | 0.105 |
Note. — The table provides equatorial coordinates (for the epoch J2000), mean -band brightness and color in quiescence, mean -band brightness and color in outburst, and color excess (Skowron et al., 2021) toward the detected objects.
OGLE-mNOVA-11 was also observed by the Neil Gehrels Swift Observatory (Gehrels et al., 2004) five times from 2023 December 6 to 30 (Appendix C). Swift’s X-Ray Telescope (XRT; Burrows et al. 2005) detected a faint X-ray source at the transient’s position, with a 0.3–10 keV count rate declining from count s-1 (December 6) to count s-1 (December 30), see Figure 2. All the data were combined to create a single X-ray spectrum, which can be fitted with a blackbody of K absorbed by a column density cm-2. If we fix to the LMC value ( cm-2), then the temperature is slightly better constrained to K. Assuming this blackbody model, the observed 0.3–10 keV luminosity is approximately erg s-1, or erg s-1 when corrected for absorption, about an order of magnitude smaller than that of ASASSN-16oh, but still substantially higher than dwarf nova X-ray luminosities.
5 Discussion and Conclusions
Overall, the optical (light curve shape, narrow emission lines, strong He ii emission) and X-ray properties of OGLE-mNOVA-11 match well those of ASASSN-16oh. We thus propose that they, and probably the remaining objects described in this study, form a homogeneous group of transient supersoft X-ray sources. We dub them ‘millinovae’ because their optical luminosities are roughly a thousand times fainter than those of ordinary classical novae. 111They should not be confused with ‘micronovae’, a much lower luminosity and much faster event seen in a number of Galactic magnetic cataclysmic variables (Scaringi et al., 2022).
Their sky location (upper panels of Figure 4) and the fact that most of the selected stars occupy a relatively narrow region in the color–magnitude diagram (middle-left panel of Figure 4) indicate that millinovae are all in the Magellanic Clouds. We used the long-term astrometric time-series of OGLE observations to measure proper motions of selected objects (Appendix D), which are consistent (with one exception) with those of stars located in the Magellanic Clouds (Gaia Collaboration et al., 2018). The radial velocities of ASASSN-16oh (Maccarone et al., 2019) and OGLE-mNOVA-11 (Appendix B) are consistent with the systemic velocities of the SMC and LMC, respectively. However, we cannot rule out that some objects are located in the foreground Milky Way disk or in background galaxies, although we consider this very unlikely.
The middle-left panel of Figure 4 shows that most objects in quiescence occupy the same region of the color–magnitude diagram, , , which indicates that the accretion disk dominates their quiescent luminosity. Such a range is consistent with the absolute magnitudes of dwarf novae in the Magellanic Clouds (Shara et al., 2003), indicating that they are essentially nova-like disks. In outburst (middle-right panel of Figure 4), objects move toward bluer colors () and higher luminosities (), as expected given the strong irradiation of the disk by the supersoft component.
The mean absolute magnitudes of millinovae in quiescence are and . In outburst, the absolute magnitudes are and , and are about a thousand times fainter than classical novae (lower-right panel of Figure 4). Millinovae are also much brighter than most known dwarf novae (lower panels of Figure 4).
The spectroscopic properties of ASASSN-16oh and OGLE-mNOVA-11, and the color evolution of OGLE-mNOVA-11 indicate that their optical outbursts likely have a disk origin. (If the optical outburst of OGLE-mNOVA-11 had a classical nova origin, the -band light curve would decline earlier than the -band; Hillman et al. 2019.) Because dwarf novae follow a well-known relation between the orbital period and absolute magnitude (both in quiescence and outburst; Patterson 2011), relatively high absolute magnitudes of millinovae indicate that their orbital periods should be on the order of a few days. Indeed, one of the analyzed objects (OGLE-mNOVA-08) shows clear eclipsing variability with a day period (Figure 5). Furthermore the rate of decline from outburst of dwarf novae, , is also known to be proportional to (the ‘Bailey relation’, Bailey 1975; see Figure 6), as expected given that longer periods imply larger disks. Most millinovae have decline rates in the range day mag-1, which corresponds to orbital periods from 3 to 15 day, assuming that the Bailey relation can be extrapolated to such long periods.
The “spreading layer” (Kippenhahn & Thomas, 1978; Inogamov & Sunyaev, 1999; Piro & Bildsten, 2004; Maccarone et al., 2019) model can naturally produce the soft X-ray temperatures that have been seen in many dwarf nova outbursts, albeit in shorter period systems such as U Gem and SS Cyg (see e.g. Wheatley et al. 2003). However, the long orbital periods of millinovae imply the existence of large disks and hence likely large mass transfer rates (e.g., Echevarría 1994) capable of producing higher X-ray luminosities. Nevertheless, there are difficulties with this model that have been noted (e.g., Kato et al. 2020) in that there are high mass transfer rate recurrent novae in quiescence (e.g. U Sco, RS Oph) where such a soft component would have been expected, but has never been seen.
The most obvious solution to explaining the properties of millinovae is then to find a way to trigger thermonuclear burning on the white dwarf surface without producing a nova flash. One such attempt to do this is the non-ejecting thermonuclear runaway event of Hillman et al. (2019), but the resulting light curves of their ASASSN-16oh model are completely different to the smooth rise and fall that is observed, not to mention the huge color variation they predict which is definitely excluded by the data.
An alternative route to such a trigger has been sought by Kato et al. (2020) who build on the already known ongoing high mass-transfer rate in these systems (their brightness and variability properties are consistent with nova-likes in the Magellanic Clouds; Shara et al. 2003), thereby keeping the disk (and white dwarf) in a hot, active state. Furthermore, they require a massive white dwarf, close to 1.2 M⊙ in order to initiate steady burning without a strong wind or shell being ejected. While we do not yet know how many of these long symmetric outburst systems produce detectable supersoft X-rays, if they were all to do so then this would have important implications for the theory of non-ejecting supersoft X-ray sources, as it would constrain the parameters of systems capable of behaving in this way. Furthermore, it would provide a route for allowing the white dwarf to continue to grow in mass, and hence become a potential Type Ia supernova progenitor.
We recognise, however, that an added complication to the Kato et al. (2020) models is that they produce close to Eddington-limited events, whereas some of our millinovae clearly do not reach such a level (Appendix E). Either there are severe geometrical constraints on the fluxes observed, or the nuclear burning does not cover the entire white dwarf surface. Nevertheless, we believe that this new group of millinovae, all likely representing long period, high mass transfer rate cataclysmic variables opens an important new route for study, with the added benefit of a well-constrained population in the Magellanic Clouds.
Acknowledgments
We thank all the OGLE observers for their contribution to the collection of the photometric data over the decades. This research was funded in part by National Science Centre, Poland, grant OPUS 2021/41/B/ST9/00252 awarded to P.M. K.L.P. acknowledges funding from the UK Space Agency. A part of this work is based on observations made with the Southern African Large Telescope (SALT), with the Large Science Programme on transients 2021-2-LSP-001 (PI: D.A.H.B.). Polish participation in SALT was funded by grant No. MEiN 2021/WK/01.
Appendix A Light Curves
Figures 7 and 8 show the light curves of the remaining stars showing long-duration symmetrical outbursts, which we classified as ‘millinovae’.
Appendix B SALT Observations of OGLE-mNOVA-11
Optical spectra of OGLE-mNOVA-11 were obtained with the Robert Stobie Spectrograph (Kobulnicky et al., 2003; Burgh et al., 2003) on the 10-m Southern African Large Telescope (SALT; Buckley et al. 2006) in long-slit mode. Two spectra were taken each night on 2023 November 22 and December 2, 3, 6, 8 and 10 with an exposure time of 1200 s for each spectrum. Observations were made with the PG0700 grating at an angle of with a slit width of , giving a resolving power of . The data were pre-reduced using the PySALT package (Crawford et al., 2010), which includes bias subtraction, gain and cross-talk corrections, and mosaicing. An Argon lamp was utilized for wavelength calibrations and the spectrophotometric standard star, EG 21, was used for flux calibration. The IRAF longslit package was used to perform the wavelength and flux calibrations, cosmic ray removal and spectra extractions. The spectra were stacked so as to obtain one average spectrum per night.
The software package, Fityk (https://fityk.nieto.pl/) (Wojdyr, 2010), was used for spectral analysis. The SALT spectra were normalized before the emission lines were modelled using the method of Shafter (1983) in order to determine the central wavelengths of the He ii (4686 Å), , and emission. This entailed fitting a Gaussian function through 95% of the line wings. In Fityk, the errors of the fitted parameters were obtained for a confidence level of (68.27%). The measured central wavelengths were then utilized to calculate the barycentric-corrected radial velocities.
Table 2 presents the barycentric radial velocities of He ii (4686 Å), H, and H. We also calculated the weighted mean radial velocity from He ii and H (the H line was detected at low significance, hence, the large uncertainties of its radial velocities). The measured values vary from to km s-1, with a mean of km s-1, which is very close to the established LMC line-of-sight velocity ( km s-1; van der Marel et al. 2002). The number of radial velocity measurements prevents us from determining the orbital period unequivocally, but it is likely to be shorter than 4 day. If we restrict the period to be longer than 1 day and assume a circular orbit, the best-fit periods are 2.42, 1.68, 2.08, and 1.89 day.
| Time | BJD | He ii | H | H | Mean RV |
| (km s-1) | (km s-1) | (km s-1) | (km s-1) | ||
| 2023 Nov 22.93 | 2460271.43 | ||||
| 2023 Dec 2.88 | 2460281.38 | ||||
| 2023 Dec 3.84 | 2460282.34 | ||||
| 2023 Dec 6.84 | 2460285.34 | ||||
| 2023 Dec 8.83 | 2460287.33 | ||||
| 2023 Dec 10.83 | 2460289.33 |
Appendix C Swift Observations of OGLE-mNOVA-11
C.1 Swift X-ray data
The Neil Gehrels Swift Observatory (Gehrels et al., 2004) performed five Target of Opportunity observations of OGLE-mNOVA-11 (Target ID 16408) between 2023 December 6 and December 30, collecting ks during each pointing, with the exception of December 17, when only ks was obtained. The X-ray Telescope (XRT; Burrows et al. 2005) was operated in Photon Counting mode, and the data were processed using HEASoft 6.32, together with the most up-to-date calibration files available. Given the faintness of the source, a small extraction region of 10 pixels () radius was used for the source, and a larger, 60 pixel () radius circle, offset from, but close to, the source was used to estimate the background contribution.
In observations where the source was detected at the 3-sigma level, the 0.3–10 keV count rate (corrected for all PSF losses) is given in Table 3. In the shorter exposure on December 17, the source was not detected, and a upper limit is given instead.
There is no significant change in the hardness ratio during these observations, though it should be noted that we are dealing with a limited number of source counts ( in total across all five observations, the vast majority below 1 keV). Therefore, all the data were combined to create a single spectrum. Cash statistics (Cash, 1979) were used to fit this spectrum with a combination of an optically-thick blackbody for the soft emission, and an optically-thin thermal plasma to account for the few harder energy photons, both absorbed by the same hydrogen column. While the optically-thin plasma can only be constrained to have a temperature of 4.2 keV ( K), the best-fitting blackbody temperature is 36 eV (418,000 K), absorbed by a column density of cm-2. This fit results in a C-stat value of 19.9 for 25 degrees of freedom.
If is fixed to the LMC value of cm-2 (HI4PI Collaboration et al., 2016), the blackbody temperature becomes slightly better constrained, with eV (temperature 607,000 K), for C-stat/dof = 21.2/26. This fit leads to an observed luminosity (in the energy range 0.3–10 keV, and at a distance of 49.59 kpc; Pietrzyński et al. 2019) of erg s-1, or erg s-1 unabsorbed.
| Obs. ID | Start time | Exp. time | Count rate |
| (UT) | (s) | ( s-1) | |
| 00016408001 | 2023-12-06 16:04:57 | 1594 | |
| 00016408002 | 2023-12-13 09:21:57 | 1499 | |
| 00016408003 | 2023-12-17 07:10:56 | 479 | |
| 00016408004 | 2023-12-20 04:27:56 | 1605 | |
| 00016408005 | 2023-12-30 16:46:56 | 1529 |
C.2 Swift UV data
OGLE-mNOVA-11 was also observed by the UltraViolet and Optical Telescope (UVOT; Roming et al. 2005) onboard the Neil Gehrels Swift Observatory. Observations took place during a period when one of the three onboard gyroscopes malfunctioned, resulting in increased noise and affecting the image quality of UVOT data (Cenko, 2023). Moreover, the nearest source detected in the UVOT images was slightly offset from the target position by and was likely a blend of the target and a nearby () bright (, ) constant star. We measured the brightness of the blend using the uvotsource tool, which performs aperture photometry. We used an aperture with a radius of centered on the position of OGLE-mNOVA-11.
We then estimated the brightness of the neighbor from the analysis of the UV–optical color–magnitude diagrams, assuming that the neighbor’s color is consistent with colors of other main-sequence stars of similar magnitude. We, therefore, created UV–optical color–magnitude diagrams by cross-matching the UVOT catalogs with the OGLE photometric map. We then calculated the mean color of main-sequence stars in the brightness range , where mag. We found:
(We also tested and and got similar results.) That enabled us to estimate the brightness of the neighbor in the UV filters:
and estimate the UV brightness of the target by subtracting the brightness of the neighbor star. Results are reported in Table 4. All magnitudes are in the Vega system.
| Obs. ID | |||
| 00016408001 | … | … | |
| 00016408002 | |||
| 00016408003 | … | … | |
| 00016408004 | … | … | |
| 00016408005 | … | … |
Appendix D OGLE Proper Motions
We calculated the proper motions of detected stars using the OGLE time-series astrometric data (OGLE-Uranus; A. Udalski et al. 2024, in preparation). In short, we measured each star’s position in individual OGLE images collected between 2010 and 2024 (in some cases 2010–2020 if later data were unavailable). We then cross-matched the positions of bright stars with those calculated using the Gaia Data Release 3 (DR3) data (Gaia Collaboration et al., 2016, 2023), taking into account their proper motion and heliocentric parallax. That allowed us to find the transformation between the positions in individual OGLE images to the equatorial coordinates in the Gaia reference frame (Lindegren et al., 2021). The proper motions in right ascension and declination for all detected stars are reported in Table 5. Because most of our targets are very faint (), the proper motion uncertainties are large (the median error bar is 2 mas yr-1). Nonetheless, in all cases (but one), the proper motions are consistent within the quoted errors with the Magellanic Clouds’ proper motions (Gaia Collaboration et al., 2018). The only exception is OGLE-mNOVA-15 with a total proper motion of mas yr-1, making it most likely to be a foreground object, and we therefore treat it as a candidate millinova.
| Object | (OGLE) | (OGLE) | (Gaia) | (Gaia) |
| (mas yr-1) | (mas yr-1) | (mas yr-1) | (mas yr-1) | |
| OGLE-mNOVA-01 | ||||
| OGLE-mNOVA-02 | ||||
| OGLE-mNOVA-03 | ||||
| OGLE-mNOVA-04 | ||||
| OGLE-mNOVA-05 | ||||
| OGLE-mNOVA-06 | ||||
| OGLE-mNOVA-07 | ||||
| OGLE-mNOVA-08 | ||||
| OGLE-mNOVA-09 | ||||
| OGLE-mNOVA-10 | ||||
| OGLE-mNOVA-11 | ||||
| OGLE-mNOVA-12 | ||||
| OGLE-mNOVA-13 | ||||
| OGLE-mNOVA-14 | ||||
| OGLE-mNOVA-15 | ||||
| OGLE-mNOVA-16 | ||||
| OGLE-mNOVA-17 | ||||
| OGLE-mNOVA-18 | ||||
| OGLE-mNOVA-19 | ||||
| OGLE-mNOVA-20 | ||||
| OGLE-mNOVA-21 | ||||
| OGLE-mNOVA-22 | ||||
| OGLE-mNOVA-23 | ||||
| OGLE-mNOVA-24 | ||||
| OGLE-mNOVA-25 | ||||
| OGLE-mNOVA-26 | ||||
| OGLE-mNOVA-27 | ||||
| OGLE-mNOVA-28 | ||||
| OGLE-mNOVA-29 |
Appendix E eROSITA upper limits
The first catalog of X-ray sources detected by the eROSITA telescope array (Predehl et al., 2021) aboard the Spektrum-Roentgen-Gamma (SRG; Sunyaev et al. 2021) satellite was recently published by Merloni et al. (2024). The number of exposures, and so the depth of the survey, depends on sky position. Thanks to the SRG scanning law, objects located near the ecliptic poles are observed most frequently. Coincidentally, the LMC is located near the south ecliptic pole.
We cross-matched our list with the eROSITA catalog (Merloni et al., 2024), but none of the objects from Table 1 were detected by eROSITA. The catalog of Merloni et al. (2024) contains X-ray sources located in the western Galactic hemisphere detected during the first six months of eROSITA operations (from 2019 December 12 to 2020 June 11, that is, ). The overlap with OGLE observations is relatively small because the OGLE-IV operations were suspended on 2020 March 18 () due to the COVID-19 pandemic. Nonetheless, we found that four objects (OGLE-mNOVA-06, OGLE-mNOVA-10, OGLE-mNOVA-19, and OGLE-mNOVA-24) were in outburst during the first eROSITA All-Sky Survey. All four objects were relatively faint and reached a peak -band magnitude of during that time. Assuming that the X-ray flux scales proportionally with peak -band flux, and taking OGLE-mNOVA-11 as a benchmark, the maximal expected X-ray flux is on the order of erg s-1 cm-2 in the 0.3–10 keV range. The expected flux in the eROSITA soft and medium energy bands (0.2–2.3 keV) is about 6% larger, erg s-1 cm-2. However, each object must have been scanned by eROSITA multiple times and the time-averaged flux is likely to be smaller.
We queried the eROSITA Upper Limits service (Tubín-Arenas et al., 2024) to get upper flux limits in the 0.2–2.3 keV range for all four objects discussed above. These limits are calculated at a 99.87% () one-sided confidence level assuming a power-law spectrum with a photon index and a column density cm-2. We obtained limits of erg s-1 cm-2 at the positions of OGLE-mNOVA-(06,10,19,24), respectively. We followed a procedure outlined in Tubín-Arenas et al. (2024) to estimate upper limits for the best-fitting absorbed blackbody model. Using the appropriate eROSITA calibration files and XSPEC version 12.14.0 (Arnaud, 1996), we estimated the energy-to-count conversion factor of cm2 erg-1, which is only 3.4% smaller than that calculated using the fiducial model in Tubín-Arenas et al. (2024). Therefore, the upper limits for our best-fitting model are only 3.4% larger than default ones. In the 0.2–5 keV range, the corresponding limits are erg s-1 cm-2. We thus conclude that the expected time-averaged flux is close to or below the current eROSITA limits. It is therefore possible that the selected stars may be detected in future eROSITA data releases once more data are processed, or may be detected in data from individual scans.